OpenStudy (pphalke):

Find the simple interest on $25000 for six months at a rate of 8.5% per year

OpenStudy (welshfella):

Interest = P*R*T ------ 100 where P = starting amount , R is the rate per cent and T = time in years.

OpenStudy (perfecta):

Convert 6 months into years. 6 months =612 years=0.5 years Find an interest by using the formula I=Pit, where I is interest, P is total principal, i is rate of interest per year, and t is total time in years. In this example P = $25000, i = 8.5% and t = 0.5 years, so I = P.i.t I = 25000.0.085.05 I = 1062.5 Find an amount by using the formula A=P+I. Since P = $25000 and I = $1062.5 we have